OpenStudy (anonymous):

Suppose that sin θ = 1/3 and 0 < θ < π/2. (a) Compute cos θ. (b) Compute sin 2θ. Hint: sin 2θ = sin(θ + θ)

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle \(\theta\) with \(\sin(\theta)=\frac{1}{3}\)

OpenStudy (anonymous):

find the other side via pyhagoras then you can take any trig ratio you like